OpenStudy (anonymous):

please help this ones pretty tough... The list price of a book is $15. But if more than 12 copies are purchased, then the price per copy is reduced by $0.20 for every copy above 12. This is, $14.80 per copy for 13 copies, $14.60 per copy for 14 copies, and so on. Which of the following is the rule for this function p(x)?

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

C.

OpenStudy (jdoe0001):

is much better to show it in 1 graphic, than scattered over 4 or so pictures

OpenStudy (anonymous):

D.

OpenStudy (anonymous):

it was impossible to save them all together sorry for the inconvenience

OpenStudy (jdoe0001):

so, let's take a look from 1 - 12 books, you pay regular retail price \(\large $15\ if \ 1\le x \le 12\) for the 13th book and above, you pay .20 less per unit so 13-12 1 unit, 0.20 less so 14-12 2 units, 0.40 less so 15-12 3 units, 0.60 less so \(\color{blue}{15 - 0.20\times unit}\) for units above the 12th one so, which one do you think?

OpenStudy (anonymous):

either b or c?

OpenStudy (jdoe0001):

b or c, so, which one would be? <rolling drums> <profuse sweat and anxiety> <biting nails>

OpenStudy (anonymous):

oh actually c?

OpenStudy (jdoe0001):

what's wrong with b?

OpenStudy (anonymous):

your adding 15+.2 instead of subtracting

OpenStudy (jdoe0001):

yep, it's c then :)

OpenStudy (anonymous):

thanks:)

OpenStudy (jdoe0001):

yw